# Personalized MR-Informed Diffusion Models for 3D PET Image Reconstruction

### v1

@@ II-D PET reconstruction with diffusion models — PET-DDS
$\mathbf{x}^{(s)}_{t_{k}}\leftarrow s_{\theta}(\mathbf{x}_{t_{k+1}},t_{k+1})$triangleright Apply $s_{\theta}$ along 2D slices
$\mathbf{x}^{0}_{t_{k+1}}\leftarrow$ Tweedie$(\mathbf{x}^{(s)}_{t_{k}},t_{k})$triangleright Estimate $\mathbb{E}(\mathbf{x}_{0}|\mathbf{x}_{t_{k}})$
$\mathbf{x}^{0}_{t_{k+1}}\leftarrow c^{-1}\cdot\mathbf{x}^{0}_{t_{k+1}}$triangleright Scale according to measurements
@@
$\qquad-\lambda^{\text{DDS}}||\mathbf{x}^{i}_{t_{k+1}}-\mathbf{x}^{0}_{t_{k+1}}||^{2}_{2}\bigr{)}$triangleright Define sub-objective
@@
$\mathbf{x}^{p}_{t_{k+1}}\leftarrow c\cdot\mathbf{x}^{p}_{t_{k+1}}$triangleright Normalize image to diffusion scale
$\mathbf{x}_{t_{k}}\leftarrow$ Re-noise$(\mathbf{x}^{(s)}_{t_{k}},\mathbf{x}^{p}_{t_{k+1}},\mathcal{N}(0,\mathrm{I}))$triangleright Re-noising step
@@ II-E Deep-learned image registration
Image registration is the problem of finding an optimal spatial transformation from a “moving” image $\mathbf{x}$ to a “fixed” image $\mathbf{f}$. For affine registration, we seek the affine map $A$ that minimizes the loss $\mathcal{L}(\mathbf{f},A\mathbf{x})$. More generally, we seek a nonlinear deformation field $\phi$ that minimizes $\mathcal{L}(\mathbf{f},\mathbf{x}\circ\phi)$. VoxelMorph [24] is a deep learning framework that parameterizes the deformation field by its inputs $\mathbf{f}$ and $\mathbf{x}$, such that we seek to learn $\phi(\cdot,\cdot)$ given a training dataset. The deformation function $\phi(\cdot,\cdot)$ is based on a U-Net architecture, and outputs the deformation field between inputs $\mathbf{f}$ and $\mathbf{x}$ (i.e. to transform a vector $\mathbf{z}$ from the space defined by $\mathbf{x}$ to the space defined by $\mathbf{f}$, we compute $\phi(\mathbf{f},\mathbf{x})(\mathbf{z})$). Its training loss sums a voxel-wise registration error $||\mathbf{f}-\mathbf{x}\circ\phi(\mathbf{f},\mathbf{x})||_{2}^{2}$ and a smoothness penalty on $\phi(\mathbf{f},\mathbf{x})$.
@@ III Related work
The work presented in this article builds on work by Singh et al., and also substantially extends earlier work proposed at the IEEE Medical Imaging Conference 2024 [32].
@@ IV-A Approach
We consider the case where we have a dataset of previously acquired paired PET-MR images $\left(\mathbf{x}_{i}^{\text{PET}},\mathbf{x}_{i}^{\text{MR}}\right)_{i=1}^{N}$. Our goal is to reconstruct the target’s PET image, from their measured sinogram $\mathbf{m}$ and MR image $\mathbf{x}^{\text{MR}}_{\text{target}}$. See Figure 1 for a graphical overview of our approach.
@@
For each image pair $\mathbf{x}_{i}$ in our dataset, we then perform affine registration of $\mathbf{x}_{i}^{\text{MR}}$ to a fixed MR image $\mathbf{x}^{\text{MR}}_{\text{affine}}$, using NiftyReg [33]. This yields an affine matrix $\mathbf{B}_{i}$ for each image.
We then train a VoxelMorph instance $\phi$ to minimize the loss $\ell(\mathbf{x}^{\text{MR}}_{\text{target}},\cdot)$ on the affine-registered datasets $\left(\mathbf{B}_{i}\mathbf{x}_{i}^{\text{MR}}\right)_{i=1}^{N}$.
@@
As all pseudo-PET images here belong in the same reference space, we may choose to sum multiple transformed PET images together to increase the diversity and statistical quality of our dataset. For this purpose, we (1) choose the number of images to sum $N$ such that $\mathbb{P}(N=n)\propto\frac{1}{n}$, (2) choose $N$ images randomly without replacement, (3) sample weights for each image as scalars $w_{i}\sim U[0,1]$, (4) scale the weights $(w_{i})$ such that they sum to 1, and (5) sum the weighted images. This method of randomly combining the images is heuristically proposed to balance the potential benefit of learning from less-noisy images (i.e. multiple images summed together) and a potential loss of diversity if we average too many images each time.
@@ IV-B1 Simulated [18F]FDG data
For reconstruction (and not training), hot lesions were simulated as uniform spheres with radii between 2 mm and 10 mm with an activity level of 1.5×1.5$\times$1.5 × gray matter. Lesions were simulated at MR resolution and down-sampled to PET resolution [26]. Sinograms were then computed as before, except now for a ground truth containing lesions.
@@ IV-B2 Real [18F]FDG data
To simulate low-count data, counts were sampled without replacement from the prompt sinogram such that in expectation each prompt sinogram bin had a mean count rate 40×40$\times$40 × lower than the full-count sinogram (i.e. 2.5% of full-count). Siemens’ scanner-specific algorithms were then used to compute contamination sinograms.
@@ IV-D4 PET forward operator
Each reconstruction method made use of the same ParallelProj projector [37], with span 11 compression and without positron range modeling. Further details of how the scanner geometry was modeled are given in Webber et al. [10].
@@ V-B Simulated data reconstruction
Figure 5: Error trade-off between areas of PET-MR agreement and areas of PET-MR mismatch (i.e. PET-only lesions) for different reconstruction algorithms. NRMSE in areas of PET-MR mismatch was calculated in lesion pixels only, while NRMSE in areas of PET-MR agreement was calculated across a set of $10\times 10\times 10$ patches of representative non-lesion brain (one per lesion).
Figure 6: Structural similarity trade-off between areas of PET-MR agreement and areas of PET-MR mismatch (i.e. PET-only lesions) for different reconstruction algorithms. SSIM in areas of PET-MR mismatch was calculated as the average SSIM in $10\times 10\times 10$ patches containing lesions, while SSIM in areas of PET-MR agreement was calculated as the average SSIM in a set of $10\times 10\times 10$ patches of representative non-lesion brain voxels (one per lesion).
Figure 7: Trade-off between lesion contrast recovery coefficient (CRC) and noise in background area for different reconstruction algorithms. CRC was calculated as the average CRC for each lesion, measured against a background region of white matter (uniform in ground truth), with the coefficient of variation measured in the same background region.
V-C Real data reconstruction
@@
Figure 10: Bias-variance trade-off comparison between different reconstruction methods from 2.5% real [18F]FDG data. For OSEM and Bowsher, white markers correspond to the lowest number of iterations (5) and black markers correspond to the maximum number of iterations. For PET-DDS-based methods, white markers correspond to the smallest value of $\lambda^{\text{DDS}}$, with black markers corresponding to the greatest value of $\lambda^{\text{DDS}}$, reflecting increased guidance strength from the pre-trained diffusion model.
@@ VI Discussion
Our work on simulated datasets reproduces and validates the work of Singh et al. [9] for PET-only reconstruction. For the real data case, we found that PET-DDS + MR did not offer a benefit over PET-DDS (see Figures 8 and 10). We hypothesize that this may be because learning from PET and MR images is a more complex problem than just learning from PET images, and that with insufficient training datasets, the benefits of MR information could not be fully realized. In the simulated case, the correspondence between the PET and MR images was much greater, which enabled the MR information to be better utilized.
@@
As presented, our method requires training a registration network and diffusion model for each reconstruction target, and so may be considered computationally intensive. However, the registration time could be massively alleviated by training a general-purpose registration function on many MR volumes in advance. Furthermore, a patch-based approach to diffusion model training (using small patches) would further cut down the computational burden of using this method [40], as would fine-tuning a pre-trained model.
There remain further avenues for exploring and improving this methodology. More advanced MR registration algorithms, including those incorporating segmentation-based losses [24], may yield higher-quality training data. For this study, we chose to use MR-MR registration rather than PET-PET registration, as it is most likely to be accurate due to the superior anatomical detail in MR relative to PET, and also allowed our registration approach to be both tracer and dose-invariant. However, PET-PET registration may also enable this methodology to apply to PET-only or PET-CT scanners.

### v2

@@ II-D PET reconstruction with diffusion models — PET-DDS
$\mathbf{x}^{(s)}_{t_{k}}\leftarrow s_{\theta}(\mathbf{x}_{t_{k+1}},t_{k+1})$ $\triangleright$ Apply $s_{\theta}$ along 2D slices
[…]
$\mathbf{x}^{0}_{t_{k+1}}\leftarrow c^{-1}\cdot\mathbf{x}^{0}_{t_{k+1}}$ $\triangleright$ Scale according to measurements
@@
$\qquad-\lambda^{\text{DDS}}||\mathbf{x}^{i}_{t_{k+1}}-\mathbf{x}^{0}_{t_{k+1}}||^{2}_{2}\bigr{)}$ $\triangleright$ Define sub-objective
@@
$\mathbf{x}^{p}_{t_{k+1}}\leftarrow c\cdot\mathbf{x}^{p}_{t_{k+1}}$ $\triangleright$ Normalize image to diffusion scale
[…]
@@ II-E Deep-learned image registration
Image registration is the problem of finding an optimal spatial transformation from a “moving” image $\mathbf{x}$ to a “fixed” image $\mathbf{f}$. For affine registration, we seek the affine map $A$ that minimizes the loss $\mathcal{L}(\mathbf{f},A\mathbf{x})$. More generally, we seek a nonlinear deformation field $\phi$ that minimizes $\mathcal{L}(\mathbf{f},\mathbf{x}\circ\phi)$. VoxelMorph [24] is a deep learning framework that parameterizes the deformation field by its inputs $\mathbf{f}$ and $\mathbf{x}$, such that we seek to learn $\phi(\cdot,\cdot)$ given a training dataset. The deformation function $\phi(\cdot,\cdot)$ is based on a U-Net architecture, and outputs the deformation field between inputs $\mathbf{f}$ and $\mathbf{x}$ (i.e. to transform a vector $\mathbf{z}$ from the space defined by $\mathbf{x}$ to the space defined by $\mathbf{f}$, we compute $\phi(\mathbf{f},\mathbf{x})(\mathbf{z})$). Its training loss sums a voxel-wise registration error $\mathbf{f}-\mathbf{x}\circ\phi(\mathbf{f},\mathbf{x})||_{2}^{2}$ and a smoothness penalty on $\phi(\mathbf{f},\mathbf{x})$.
@@ III Related work
The work presented in this article builds on work by Singh et al. [9], and also substantially extends earlier work proposed at the IEEE Medical Imaging Conference 2024 [32].
@@ IV-A Approach
We consider the case where we have a dataset of previously acquired paired PET-MR images $\left(\mathbf{x}_{i}^{\text{PET}},\mathbf{x}_{i}^{\text{MR}}\right)_{i=1}^{N}$. Our goal is to reconstruct the target’s PET image, from their measured sinogram $\mathbf{m}$ and MR image $\mathbf{x}^{\text{MR}}_{\text{target}}$. To this end, we seek to generate pseudo-PET images that are in the anatomical space defined by our target subject, and thereby personalize an SGM for reconstruction. Hypothetically, this personalized diffusion model has an easier learning task than a non-personalized diffusion model, as it only needs to learn PET distributions in the target anatomy (instead of generalizing to unseen brain anatomies). The model is also trained on less-noisy pseudo-PET images, that have anatomical features of relevance, so the prior information being incorporated is of greater relevance. See Figure 1 for a graphical overview of our approach.
@@
For each image pair $\mathbf{x}_{i}$ in our dataset, we then perform affine registration of $\mathbf{x}_{i}^{\text{MR}}$ to the same affine space (defined relative to a fixed MR image $\mathbf{x}^{\text{MR}}_{\text{affine}}$, which may be the subject’s MR image but does not have to be), using NiftyReg [33]. This yields an affine matrix $\mathbf{B}_{i}$ for each image.
We then train a VoxelMorph instance $\phi$ to minimize the loss $\ell(\mathbf{B}_{\text{target}}\mathbf{x}^{\text{MR}}_{\text{target}},\cdot)$ on the affine-registered datasets $\left(\mathbf{B}_{i}\mathbf{x}_{i}^{\text{MR}}\right)_{i=1}^{N}$.
@@
As all pseudo-PET images here belong in the same reference space, we may choose to sum multiple transformed PET images together to increase the diversity and statistical quality of our dataset. For this purpose, we (1) choose the number of images to sum $N$ such that $\mathbb{P}(N=n)\propto\frac{1}{n}$, (2) choose $N$ images randomly without replacement, (3) sample weights for each image as scalars $w_{i}\sim U[0,1]$, (4) scale the weights $(w_{i})$ such that they sum to 1, and (5) sum the weighted images. This method of randomly combining the images is heuristically proposed to balance the potential benefit of learning from less-noisy images (i.e. multiple images summed together) and the potential risk that all images appear too similar (if too many images are summed each time and no images in the resulting dataset have fine-grained features).
@@ IV-B1 Simulated [18F]FDG data
For reconstruction (and not training), hot lesions were simulated as uniform spheres with radii between 2 mm and 10 mm with an activity level of $1.5\times$ gray matter. Lesions were simulated at MR resolution and down-sampled to PET resolution [26]. Sinograms were then computed as before, except now for a ground truth containing lesions.
@@ IV-B2 Real [18F]FDG data
To simulate low-count data, counts were sampled without replacement from the prompt sinogram such that in expectation each prompt sinogram bin had a mean count rate $40\times$ lower than the full-count sinogram (i.e. 2.5% of full-count). Siemens’ scanner-specific algorithms were then used to compute contamination sinograms.
@@ IV-D4 PET forward operator
Each reconstruction method made use of the same 3D ParallelProj projector [37], with span 11 compression and without positron range modeling. Further details of how the scanner geometry was modeled are given in Webber et al. [10].
@@ V-B Simulated data reconstruction
Figure 5: Error trade-off between areas of PET-MR agreement and areas of PET-MR mismatch (i.e. PET-only lesions) for different reconstruction algorithms. NRMSE in areas of PET-MR mismatch was calculated in lesion pixels only, while NRMSE in areas of PET-MR agreement was calculated across a set of $10\times 10\times 10$ patches of representative non-lesion brain (one per lesion). For OSEM and Bowsher, white markers correspond to the lowest number of iterations (5) and black markers correspond to the maximum number of iterations. For PET-DDS-based methods, white markers correspond to the smallest value of $\lambda^{\text{DDS}}$, with black markers corresponding to the greatest value of $\lambda^{\text{DDS}}$, reflecting increased guidance strength from the pre-trained diffusion model.
Figure 6: Structural similarity trade-off between areas of PET-MR agreement and areas of PET-MR mismatch (i.e. PET-only lesions) for different reconstruction algorithms. SSIM in areas of PET-MR mismatch was calculated as the average SSIM in $10\times 10\times 10$ patches containing lesions, while SSIM in areas of PET-MR agreement was calculated as the average SSIM in a set of $10\times 10\times 10$ patches of representative non-lesion brain voxels (one per lesion). See Figure 5 for marker meanings.
Figure 7: Trade-off between lesion contrast recovery coefficient (CRC) and noise in background area for different reconstruction algorithms. CRC was calculated as the average CRC for each lesion, measured against a background region of white matter (uniform in ground truth), with the coefficient of variation measured in the same background region. See Figure 5 for marker meanings.
V-C Algorithm variants
V-C1 PET-PET registration for pseudo-PET images
(a) Example reconstruction of the algorithm variants PET-reg + PET-DDS and MR-reg + PET-DDS + MR. All SGM-based methods used the same hyperparameter $\lambda^{DDS}=0.4$.
(b) Error trade-off between areas of PET-MR agreement and areas of PET-MR mismatch (i.e. PET-only lesions) for selected algorithms. (See Figure 5 for further details.)
Figure 8: Results (on simulated data at 2.5% count) for variants of the proposed method. PET-reg + PET-DDS is a variant of the proposed method that uses PET-based registration to generate pseudo-PET images. MR-reg + PET-DDS + MR is a variant that uses MR-registered pseudo-PET images, and additionally provides the subject’s MR image as a side-channel to the SGM (during training and reconstruction).
We investigated using PET-PET registration instead of MR-MR registration for producing pseudo-PET images. In Figure 8(a), we show example quantitative and qualitative results from the method ‘PET-reg + PET-DDS‘ (PET-DDS trained with PET-registered pseudo-PET images). This figure shows that reconstructions from an SGM trained on PET-registered images have fewer anatomical details present than their MR-registered equivalent. Compared to standard PET-DDS, background noise appears less well suppressed. While some structures appear clearer than in PET-DDS (e.g. the cortical folds at the top of the image), others appear more hallucinatory (e.g. the thalamus, appearing as a central inverted v-shape). These observations are bourne out by the quantitative performance of the method, summarized in Figure 8(a).
V-C2 Including MR information as an additional channel for MR-reg + PET-DDS
We also investigated supplying the subject’s MR image to the SGM during training and reconstruction. In our experiments, the MR image was supplied as a second channel to the PET image. Figure 8 shows the qualitative and quantitative results from experiments with this variant, demonstrating that there seems to be no additional benefit to including the MR-information in this secondary way.
V-D Real data reconstruction
@@
Figure 11: Bias-variance trade-off comparison between different reconstruction methods from 2.5% real [18F]FDG data. See Figure 5 for marker meanings.
V-E Computational efficiency
Table I gives wall-clock timings for each approach. Training a registration network took up to 20.0 GB of GPU memory, while training a diffusion model took up to 19.7 GB of GPU memory. Performing reconstruction with the diffusion model and model-based data fidelity updates took up to 22.5 GB of GPU memory.
TABLE I: Wall-clock timings recorded for each algorithm on reconstruction from 3D simulated FDG data.
Method Time
Pre-train Train Recon
OSEM N/A N/A 4 s
Bowsher N/A N/A 38 s
PET-DDS N/A 5 h 44m 160 s
PET-DDS+MR N/A 4h 23m 252 s
MR-reg + PET-DDS 11h 52m 4 h 27m 160 s
@@ VI Discussion
Our work on simulated datasets reproduces and validates the work of Singh et al. [9] for PET-only reconstruction. For the real data case, we found that PET-DDS + MR did not offer a benefit over PET-DDS (see Figures 9 and 11). We hypothesize that this may be because learning from PET and MR images is a more complex problem than just learning from PET images, and that with insufficient training datasets, the benefits of MR information could not be fully realized. In the simulated case, the correspondence between the PET and MR images was much greater, which enabled the MR information to be better utilized.
@@
As presented, our method requires training a registration network and diffusion model for each reconstruction target, and so may be considered computationally intensive (see Table I). However, the registration time could be massively alleviated by training a general-purpose registration function on many MR volumes in advance. Furthermore, a patch-based approach to diffusion model training (using small patches) would further cut down the computational burden of using this method [40], as would fine-tuning a pre-trained model.
Beyond improving efficiency, several avenues remain for further developing this methodology. More advanced MR registration algorithms, including those incorporating segmentation-based losses [24], may yield higher-quality training data. The pseudo-PET data generated could be used with different score-based prior approaches such as PET-DPS [9] or likelihood-scheduling for PET reconstruction [10]. For this study, we chose to use MR-MR registration rather than PET-PET registration, as it is most likely to be accurate due to the superior anatomical detail in MR relative to PET, and also allowed our registration approach to be both tracer and dose-invariant. Our initial results with PET-PET registration suggest that it is not suitable for ultra-low dose imaging; however, it remains to be seen whether results improve with higher dose levels.